MAT 141, HONORS CALCULUS, SUNY AT STONY BROOK, FALL 1999

Lecturer: Prof. Christopher Bishop, bishop@math.sunysb.edu, office is 4-112 in Math Building. Office hours are MF 11:30-12:30.

Text: Calculus, 9th edition, Thomas and Finney.

Syllabus: We will cover approximately Chapters 1, 2, 3, 4 and part of Chapter 5 with certain omissions and additions of sections. A more detailed syllabus will be provided.

Classdays. Classes begin on Wednesday, September 1 and end on Monday, December 13.

Holidays: There are no classes on Labor Day (Monday, September 6), Yom Kipur (Monday, September 20) and Thanksgiving (November 25-26). Wednesday, September 8 will follow a Monday schedule, so we will have lecture on that day.

Lecture times: Monday and Friday, 12:40-2:00 pm.

Lecture location: The lectures will be on Mondays and Fridays, 12:40-2:00 in room 152 of Light Engineering.

Recitation times and locations: There are two recitiations.

• Section 1 meets Tuesday-Thursday at 11:20-12:15 in ESS 183 (Earth and Space Sciences). The TA is Gregory Janks. His office number is 2-115 in the Math Building and his email is gbj@math.subysb.edu.

• Section 2 meets Tuesday-Thursday at 12:50-1:45 in room P-128 in Physics building. The TA is Peter Hennes. His office number is 2-121 in the Math

Building and his email is phennes@math.sunysb.edu.

Homework. A list of problems from the text and supplementary material will be handed out. The recitation instructor will assign, collect and return homework. They may also give weekly quizzes based on the assigned problems.

Examinations. There will be two midterm tests: Friday, October 15 and Friday, November 19. These will be given during the regular lecture time. The final exam is on Friday, December 17 at noon to 3pm. The location will be announced in class during the last week of the semester.

Early Midterm: There will be an evening exam at 8:30 pm on Tuesday, September 14 in Javits 103. This is part of the placement process and is given to all students in calculus classes. More information and a practice exam can

be found on the web at http://www.math.sunysb.edu/earlyexam

Grades. Grades are computed using: Recitation 10%, Early Midterm 10%, Midterms 20% each, Final 40%.

Disabilities. If you have any condition, such as physical or mental disability, that will make it difficult to complete the course or that will require extra time on examinations, notify your lecturer during the first two weeks of class so that appropriate arrangements can be made.

Office Hours. The lecturer and TAs will announce office hours in class. Math Learning Center: (Physics A-125/127) This is located on the

A level of the physics building and provides help with pre-calculus material. The same room also contains the Calculus Resource Room where you can go for help with Calculus classes. It is staffed most days and some evenings; an exact schedule will be posted in the room. Note that this room is closed during final exam week. It also gets crowded before exams, so don’t wait until the last minute to get help there.

Calculus Online: There are many math sites on the internet. For information on the history of Calculus and short biographies of famous mathe-

maticians, try http://www-groups.dcs.st-and.ac.uk/~history and click

on History Topics or the Biography Index. You can find information about

math classes at Stony Brook at www.math.sunysb.edu \end{verbatim}

Drop dates: Last day for Add/Drop is Wednesday, September 15. Last day to switch to lower numbered math course is Friday, October 22. The last day to withdraw from a course is Tuesday, November 23.